# Oscillons from $Q$-balls

### v1

@@ II Renormalization Group Perturbation Expansion
  • Solve RGEs and set the renormalization scale in such a way that all secular terms in the expansion are removed, i.e. choose a subtraction scheme. This gives a renormalized solution that is a global approximation of the true solution valid in some range of amplitudes.
@@
The second term on the right-hand side is a resonant term as it belongs to the kernel of the operator $\bigl{(}\partial_{t}^{2}+1\bigr{)}$. Naively, it seems that we obtain an infinite contribution to the perturbation series, i.e. $\bigl{(}\partial_{t}^{2}+1\bigr{)}^{-1}\mathrm{e}^{\mathrm{i}t}=$$\infty$”. Thus, we are in a situation somewhat analogous to what happens at a first loop correction in most QFT calculations. In order to make progress, we need to invoke regularization.
@@ III.1 A generic case
where $\partial^{2}\equiv\partial_{t}^{2}-\partial_{x}^{2}$, a naive, small-amplitude perturbation expansion, i.e.
@@
and where $\mathcal{S}(\theta,\bar{\theta})$ is the secular term, i.e. a solution to the equation
@@
Going with the minimal choice, $\mathcal{S}_{0}=\mathcal{S}$, the only way how to make the RG equation independent of the form of the secular term amounts to taking a particular combination of partial derivatives, i.e.
@@ III.2 Some non-generic cases
The renormalized solution (III.1) was derived under the tacit assumption that at least one of the coefficients $a_{3}$ and $a_{4}$ is non-zero. Furthermore, if the coefficients saturate the condition (41), i.e. $010a_{3}^{2}+9a_{4}=$ 0, the RGPE algorithm is also valid, (or, indeed, mute) as the resonant term disappears and we need to proceed to higher orders in the perturbation series.
Let us now comment on such non-generic situations. To begin with, let us execute the RGPE algorithm for potentials that have $a_{3}=a_{4}=0$, i.e.
@@ V.1 Integrable complex sine-Gordon
The sense, in which this equation is “close” to RG equation (37) is that it has exactly the same $Q$-ball solution given in Eq. (39). Furthermore, the Lagrangian describing $\mathbb{C}$sG model differs from (38) by a multiplicative factor of the form $1+\mathcal{O}\bigl{(}|\Psi|^{2}\bigr{)}$, i.e.
@@
For our purposes and for simplicity, let us consider the coincident, stationary two Q-ball solutions, i.e. $v_{1,2}=a_{1,2}=0$, describing a bound state of two $Q$-balls located on top of each other and centered at the origin with zero total momentum.
@@ V.2 Modulations in $\phi^{3}$ oscillons
We consider the simplest model belonging to the generic universality class i.e., $\phi^{3}$ theory. As we previously verified the single $Q$-ball solution was not able to capture the modulated oscillons. Now, we use the two $Q$-ball state and insert it into the renormalized solution
@@ VI Corrections to RG perturbation expansions
As an example, let us consider the sine-Gordon model, i.e. $a_{3}=a_{5}=0$, $a_{4}=1/6$, and $a_{6}=-1/120$. Plugging in the corresponding $Q$-ball solution
@@
As another example, let us consider a cubic model, i.e. $a_{3}=1$ and $a_{4,5,6}=0$. The corresponding $Q$-ball solution reads
@@ VII Summary
In the present work, we have resolved the long-standing, important problems related to oscillons.
@@
In the generic universality class, defined by the complex $\phi^{4}$ equation, we have models with the cubic and/or quartic self-interaction terms in the potential. Higher terms may also exist. Oscillons of the $\phi^{3}$, the reverse $\phi^{4}$, and the double well $\phi^{4}$ and many other Pujo-2 belong to this class. Here the same single $Q$-ball solution generates an approximation which very well reproduces dynamics of the simplest, unexcited i.e., unmodulated oscillons. This RG-based approximation is similar, but not identical to the standard FFHL expansion Fodor:2008es .

### v2

@@ II Renormalization Group Perturbation Expansion
  • Solve RGEs and set the renormalization scale in such a way that all secular terms in the expansion are removed, i.e., choose a subtraction scheme. This gives a renormalized solution that is a global approximation of the true solution valid in some range of amplitudes.
@@
The second term on the right-hand side is a resonant term as it belongs to the kernel of the operator $\bigl{(}\partial_{t}^{2}+1\bigr{)}$. Naively, it seems that we obtain an infinite contribution to the perturbation series, i.e., $\bigl{(}\partial_{t}^{2}+1\bigr{)}^{-1}\mathrm{e}^{\mathrm{i}t}=$$\infty$”. Thus, we are in a situation somewhat analogous to what happens at a first loop correction in most QFT calculations. In order to make progress, we need to invoke regularization.
@@ III.1 A generic case
where $\partial^{2}\equiv\partial_{t}^{2}-\partial_{x}^{2}$, a naive, small-amplitude perturbation expansion, i.e.,
@@
and where $\mathcal{S}(\theta,\bar{\theta})$ is the secular term, i.e., a solution to the equation
@@
Going with the minimal choice, $\mathcal{S}_{0}=\mathcal{S}$, the only way how to make the RG equation independent of the form of the secular term amounts to taking a particular combination of partial derivatives, i.e.,
@@ III.2 Some non-generic cases
The renormalized solution (III.1) was derived under the tacit assumption that at least one of the coefficients $a_{3}$ and $a_{4}$ is non-zero. Furthermore, if the coefficients saturate the condition (41), i.e., $010a_{3}^{2}+9a_{4}=$ 0, the RGPE algorithm is also valid, (or, indeed, mute) as the resonant term disappears and we need to proceed to higher orders in the perturbation series.
Let us now comment on such non-generic situations. To begin with, let us execute the RGPE algorithm for potentials that have $a_{3}=a_{4}=0$, i.e.,
@@ V.1 Integrable complex sine-Gordon
The sense, in which this equation is “close” to RG equation (37) is that it has exactly the same $Q$-ball solution given in Eq. (39). Furthermore, the Lagrangian describing $\mathbb{C}$sG model differs from (38) by a multiplicative factor of the form $1+\mathcal{O}\bigl{(}|\Psi|^{2}\bigr{)}$, i.e.,
@@
For our purposes and for simplicity, let us consider the coincident, stationary two Q-ball solutions, i.e., $v_{1,2}=a_{1,2}=0$, describing a bound state of two $Q$-balls located on top of each other and centered at the origin with zero total momentum.
@@ V.2 Modulations in $\phi^{3}$ oscillons
We consider the simplest model belonging to the generic universality class, i.e., $\phi^{3}$ theory. As we previously verified the single $Q$-ball solution was not able to capture the modulated oscillons. Now, we use the two $Q$-ball state and insert it into the renormalized solution
@@ VI Corrections to RG perturbation expansions
As an example, let us consider the sine-Gordon model, i.e., $a_{3}=a_{5}=0$, $a_{4}=1/6$, and $a_{6}=-1/120$. Plugging in the corresponding $Q$-ball solution
@@
As another example, let us consider a cubic model, i.e., $a_{3}=1$ and $a_{4,5,6}=0$. The corresponding $Q$-ball solution reads
@@ VII Summary
In the present work, we have resolved the long-standing, important problems related to oscillons. [^2] This work is also presented in a shortened, Letter version in Ref. Letter .
@@
In the generic universality class, defined by the complex $\phi^{4}$ equation, we have models with the cubic and/or quartic self-interaction terms in the potential. Higher terms may also exist. Oscillons of the $\phi^{3}$, the reverse $\phi^{4}$, and the double well $\phi^{4}$ and many other Pujo-2 belong to this class. Here the same single $Q$-ball solution generates an approximation which very well reproduces dynamics of the simplest, unexcited, i.e., unmodulated oscillons. This RG-based approximation is similar, but not identical to the standard FFHL expansion Fodor:2008es .